Calculate the molar mass of the following substances.
 Ethyne, C2H2

Grade:12th pass

2 Answers

Harshit Singh
2 years ago
Dear Student

Ethyne, C2H2 = 2 × 12 + 2 × 1 = 24 + 2 = 26 g

Thanks
Shraddha Singh
2 years ago
Dear Student,
 
The molar mass of ethyne C2 H= 2 × Mass of C + 2 × Mass of H
= 2 × 12 + 2 × 1 = 24 + 2 = 26 g/mol
 
Hope it helps!
Thank you ✌️